OpenStudy (anonymous):

An object is launched at 19.6 meter per second from a 58.8 meter tall platform. The equation for the objects height h at time t seconds after launch is h(t)=4.9t2+19.6t+58.8, where h is in meters. When does the object strike the ground?

OpenStudy (anonymous):

set \[4.9t^2+19.6t+58.8=0\] solve for \(t\) may be easier to solve \[49t^2+196t+588=0\]

OpenStudy (anonymous):

oh i see the problem the equation is wrong should be \[-4.9t^2+19.6t+58.8=0\]